Let X1, X2, · · · , Xn ∼ N(µ, σ2 ). Show that the MLE of µ is efficient.
Answer #1

MLE of \mu will be the sample mean = \frac{\sum_{i=1}^{n}Xi}{n}

You can find this by taking a partial dervative of the log- likelihood function with respect to \mu and the simply equating it to 0, as follows:
x constant Taking logs logL -nlogơ- 1Σ(4-1),constant Differentiating with respect to μ and σ gives i-1 Setting these to zero


Now, an estimator is said to be efficient if the mean squared error becomes 0, as n\rightarrow \infty, where n is the sample size:

Proof:

MSE(\mu) = Var(〉. Xi/n) + (Bias(\sum Xi/n))^{2}


Now, as this is an unbiased estimator, it means that Bias=0

MSE =

Var(〉. Xi/n)


={\frac{1}{n^{2}}}{Var({X1+X2...Xn})}

=  {\frac{1}{n^{2}}}{(Var({X1)+Var(X2)...Var(Xn)})}

= {\frac{1}{n^{2}}}{n* Var(X)}

=\frac{1}{n}{ Var(X)}

=\frac{1}{n}{\sigma^{2}}


Now as n\rightarrow \infty

\frac{1}{n}{\sigma^{2}}\rightarrow 0

Thus, MLE of\Mu\Mu\mu is efficient

=